OpenStudy (anonymous):

tan(x/3)=sqt3 solution set

OpenStudy (anonymous):

tan(x/3)=sqt3 x/3 = arc tan sqrt(3) x = (arc tan sqrt (3)) * 3 x = 180

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome, if it help u give me medal or become my fans :D

OpenStudy (anonymous):

haha will do im new to this site...